OpenStudy (anonymous):

Cookie ate three fourths dozen of choclate chip cookies in one bite. How many cookies did Cookie eat in that bite? Please explain and leave the answer please and thanks! :(

OpenStudy (superflytyguy):

Dozen = 12 4/4 = 12 Because 4 / 4 = 1. 2/4 or 1/2 = 6 cookies 3/4 = .75 Multiply .75 times 12 cookies

OpenStudy (anonymous):

THX :)

OpenStudy (superflytyguy):

You're welcome! :)